OpenStudy (anonymous):

Find the distance between the points given. Simplify all irrational answers. (0, 6) and (5, 12) Please answer and then explain how u did it much appreciated : )

jimthompson5910 (jim_thompson5910):

d=sqrt((x2-x1)^2+(y2-y1)^2) d=sqrt((5-0)^2+(12-6)^2) d=sqrt(5^2+6^2) d=sqrt(25+36) d=sqrt(61) So the exact distance between the two points is \[\large \sqrt{61}\] units